# KAGNNs: Kolmogorov-Arnold Networks meet Graph Learning

### v3

@@ 3 KAN-based GNN Layers
We next derive variants of the GIN and GCN models which use KANs to transform the node features instead of fully-connected layers or MLPs.
@@ 3.1 The KAGIN Layer
With theoretically-sound KANs (i. e., with arbitrarily complex components), this architecture is exactly as expressive as the vanilla GIN model with infinite layer width. In practice, we employ two different families of functions to serve as the base components, leading to two different layers: BS-KAGIN, based on B-splines, as in the original paper; and RBF-KAGIN, based on radial basis functions (RBFs), which were proposed as a more computationally-efficient alternative. While these do not guarantee universal approximation, the empirical results in the original paper demonstrate the great expressive power of KANs [12], motivating our empirical study.
@@ 3.2 The KAGCN Layer
GCN-based architectures have achieved great success in node classification tasks. While in our experiments we evaluate KAGIN on node classification datasets, the objective advantage of GCN over GIN on some of the datasets does not facilitate a fair estimation of KANs’ potential in this context. To this end, we also propose a variant of the GCN model. Specifically, we substitute the parameters and ReLU function of the standard GCN [15] model with a single KAN layer (defined in Equation 6) to obtain the KAGCN layer:
@@ 4 Empirical Evaluation
In this section, we compare both KAGIN and KAGCN models against the GIN and GCN models in the following tasks: node classification, graph classification, and graph regression. The code for reproducing the results is available at anonymized. All models are implemented with PyTorch [32]. For KAN layers, we rely on publicly available implementations for B-splines  [^1] https://github.com/Blealtan/efficient-kan and for RBF  [^2] https://github.com/ZiyaoLi/fast-kan. For a given task, all models have consistent architectures, differing as little as possible.
@@ 4.1 Node classification
To evaluate the performance of GNNs with KAN layers in the context of node classification, we use $7$ well-known datasets of varying sizes and types, including homophilic (Cora, Citeseer [15] and Ogbn-arxiv [33]) and heterophilic (Cornell, Texas, Wisconsin, Actor) networks. The homophilic networks are already split into training, validation, and test sets, while the heterophilic datasets are accompanied by fixed $10$-fold cross-validation indices.
@@
We perform two types of experiments. First, we compare our KAGCN model on the homophilic datasets to the GCN proposed in the state of the art approach in [34]. The results are shown in Table 1. One can observe that the classical GCN outperforms the KAN-based model in 2 out of 3 datasets, albeit with a narrow difference and confidence overlaps in all of the comparisons. Note that current setting is built to favor the classical GNNs, as it relies on techniques such as dropout, jumping knowledge, batch normalization and preliminary linear transformations. Hence, although the MLP-based GCN enjoys a plethora of assistive techniques examined in the past decades, the KAN-based counterpart exhibits a close performance. We should note here that these assistive techniques have not been developed yet for KAN-based architectures, and our results aim at motivating the development of such techniques.
Table 1: Average classification accuracy ($\pm$ standard deviation) and model parameters based on Luo et al.[34] using Fastkan, KAGCN, GCN on the Cora, CiteSeer, and Arxiv datasets.
@@
Table 2: Hyperparameter ranges for node classification
Model Learning Rate Hidden Layers Hidden Dim. Grid Size Spl. Order
GIN $\left[1e-5,1e-2\right]$ $\left[1,4\right]$ $\left[8,1024\right]$ NA NA
BS-KAGIN/KAGCN $\left[1e-5,1e-2\right]$ $\left[1,4\right]$ $\left[2,64\right]$ $\left[1,8\right]$ $\left[1,4\right]$
RBF-KAGIN/KAGCN $\left[1e-5,1e-2\right]$ $\left[1,4\right]$ $\left[2,64\right]$ $\left[2,32\right]$ NA
Our second approach compares the vanilla versions of the neural architectures when we replace the MLP with KAN, as delineated in sections 3. We only use message passing layers, skip connections and an output layer. We do not utilize dropout, batchnorm, jumping knowledge etc. The reason we constrain ourselves to the bare minimum is to isolate and clarify the effect of KAN on the model’s performance. In other words, we aspire to compare the classical MLP-based GNNs as if they were in the same nascent stage as their KAN-based counterparts. Having said that, we perform an auxiliary experiment afterwards, to position the KAGCN in the SOTA and clarify its current utility.
We then evaluate the $10$ models on the test set and report the average accuracy. For the heterophilic datasets, we report the average accuracy across the $10$ folds. In order for all models to exploit the same quantity of information, we fix the number of message-passing layer dataset-wise, with $2$ for Cora and CiteSeer, $3$ for Texas, Cornell, Wisconsin, and Ogbn-arxiv and $4$ for Actor.
For every dataset and model, we tune the values of the hyperparameters using the Optuna package [35]. For each parameterization, we train $5$ models and select the parameterization that yields the lowest mean validation loss (cross-entropy) across all $5$ models. We set the number of iterations of Optuna equal to $100$ trials with a TPE Sampler. We use early stopping with a patience of $50$. Once the best hyperparameter values are found, we evaluate the models on the test set. For the homophilic networks, we initialize and train $10$ different models with the best hyperparameters (using $10$ different random seeds). The ranges for hyperparameters are different for each model. Indeed, for the same number of ”neurons”, a KAN-based layer will have a much larger number of parameters than an MLP-based layer (see Appendix 5.2 for examples). In order to avoid giving an unfair advantage to KAN-based layers, we thus allowed the MLP layers to be much larger, so that the largest possible model in all cases had around the same number of parameters. The hyperparameters ranges are given in table 2.
Table 3: Average classification accuracy ($\pm$ standard deviation) of the KAGIN, GIN, KAGCN, and GCN models on the $7$ node classification datasets.
@@
GIN 67.48$\pm$6.53 61.04$\pm$6.42 35.68$\pm$12.9 41.70$\pm$1.70 45.14$\pm$6.32 43.24$\pm$2.68 26.43$\pm$0.86
BS-KAGIN 76.28$\pm$0.98 63.42$\pm$2.06 30.53$\pm$5.59 48.92$\pm$1.08 62.30$\pm$0.23 50.78$\pm$1.54 28.00$\pm$0.37
RBF-KAGIN 72.34$\pm$3.74 44.26$\pm$12.0 21.77$\pm$13.2 47.49$\pm$1.85 59.92$\pm$1.88 51.73$\pm$1.51 28.91$\pm$0.40
GCN 75.00$\pm$4.11 62.35$\pm$3.87 40.60$\pm$6.75 41.27$\pm$1.96 55.32$\pm$3.27 52.02$\pm$1.12 28.64$\pm$0.24
BS-KAGCN 78.99$\pm$0.80 66.81$\pm$1.25 52.67$\pm$4.07 48.73$\pm$1.58 58.70$\pm$0.52 56.18$\pm$1.09 27.07$\pm$0.27
RBF-KAGCN 63.59$\pm$7.10 48.00$\pm$8.32 54.56$\pm$3.99 45.86$\pm$2.10 58.43$\pm$1.21 53.35$\pm$1.38 27.10$\pm$0.31
@@ Results.
The results are given in Table 3. We see that KAN-based variants beat MLP-based models on all datasets, while the B-splines-based variant outperforms both other models on $5/7$ datasets. We also observe that RBF-KAGIN seems to exhibit a high instability, especially on Citeseer and Ogbn-arxiv, where its performance is much lower than that of other models, and its variance a lot higher. We notice that the simpler GCN-based models are more efficient than the more complex GIN-based models on all three homophilic datasets, while the opposite is true on heterophilic datasets. This is coherent with the inherent complexity discrepency between both families of datasets [36].
Overall, B-splines are more effective for both graph learning models, the difference being clear mostly in homophilic networks. Overall we contend that KAN has a positive impact on both, the GIN and the GCN architecture. Moreover, the introduction of KAN does not alleviate the inherent disadvantage of the model e. g., GCN’s low accuracy on heterophilic networks, however, it improves the models in the majority of the datasets.
@@ 4.2 Graph Classification
In this set of experiments, we compare the KAGIN models against GIN on standard graph classification benchmark datasets [37]. We experiment with the following $7$ datasets: (1) MUTAG, (2) DD, (3) NCI1, (4) PROTEINS, (5) ENZYMES, (6) IMDB-B, (7) IMDB-M. The first $5$ datasets come from bio- and chemo-informatics, while the last $2$ are social interaction datasets.
Experimental setup. We follow the experimental protocol proposed in [38]. We kept our architecture simple. First, a number of message-passing layers extract node representation, whose model is either an MLP (resp. a KAN). We then use sum pooling to obtain graph representations. Finally, an MLP (resp. a KAN) with same architecture as those found in the message-passing layer, is used as a classifier. A softmax is used for predicting class probabilities. Thus, we perform $10$-fold cross-validation, while within each fold a model is selected based on a $90\%/10\%$ split of the training set. We use the pre-defined splits provided in [38]. We use the Optuna package to select the model that achieves the lowest validation cross-entropy loss over $100$ iterations. For a fair comparison, we set the number of message-passing layers of all architectures to a fixed value for each dataset. On MUTAG, PROTEINS, IMDB-B and IMDB-M, we set the number of layers to $2$. On DD, we set it to $3$. On ENZYMES, we set it to $4$ and finally, on NCI1, we set it to $5$. We train each model for $1,000$ epochs (early stopping with a patience of $20$ epochs) by minimizing the cross entropy loss. We use the Adam optimizer for model training [39]. We apply batch normalization [40] and dropout [41] to the output of each message-passing layer. The ranges for hyperparameters search are given in table 4. Once the best hyperparameters are found for a given split, we train $3$ different models on the training set of the split and evaluate them on the test set of the split. This yields $3$ test accuracies for this split, and we compute their average to obtain the performance for this split.
Table 4: Hyperparameter ranges for node classification
Table 5: Average classification accuracy ($\pm$ standard deviation) of the KAGIN and GIN models on the $7$ graph classification datasets.
@@
GIN 85.26 $\pm$ 7.30 74.78 $\pm$ 4.65 78.32 $\pm$ 2.49 71.01 $\pm$ 2.81 53.28 $\pm$ 1.23 73.13 $\pm$ 3.97 50.27 $\pm$ 4.01
BS-KAGIN 85.08 $\pm$ 6.07 75.37 $\pm$ 3.37 79.44 $\pm$ 1.67 74.09 $\pm$ 3.60 42.72 $\pm$ 8.94 73.37 $\pm$ 4.03 49.91 $\pm$ 3.47
RBF-KAGIN 85.48 $\pm$ 5.36 75.66 $\pm$ 3.08 78.33 $\pm$ 1.81 73.29 $\pm$ 3.56 41.17 $\pm$ 8.32 71.97 $\pm$ 3.81 48.76 $\pm$ 4.53
GCN 56.34 $\pm$ 15.1 63.24 $\pm$ 3.82 60.42 $\pm$ 1.01 60.46 $\pm$ 2.06 22.06 $\pm$ 5.58 68.70 $\pm$ 3.63 46.24 $\pm$ 4.83
BS-KAGCN 75.55 $\pm$ 8.14 73.31 $\pm$ 4.45 71.59 $\pm$ 3.31 75.55 $\pm$ 3.62 57.67 $\pm$ 8.03 70.83 $\pm$ 3.95 50.00 $\pm$ 4.41
RBF-KAGCN 73.52 $\pm$ 11.6 68.30 $\pm$ 4.53 63.95 $\pm$ 7.80 74.18 $\pm$ 3.64 59.39 $\pm$ 8.03 72.60 $\pm$ 4.61 48.53 $\pm$ 4.03

### v4

@@ 4 KAN-based GNN Layers
We next derive three KAGNN layers, i.e. variants of the GIN, GCN and GAT models which use KANs to transform the node features instead of fully-connected layers or MLPs.
@@ 4.1 The KAGIN Layer
With theoretically-sound KANs (i. e., with arbitrarily complex components), this architecture is exactly as expressive as the vanilla GIN model with infinite layer width. In practice, we employ two different families of functions to serve as the base components, leading to two different layers: BS-KAGIN, based on B-splines, as in the original paper; and RBF-KAGIN, based on radial basis functions (RBFs), which were proposed as a more computationally-efficient alternative. While these do not guarantee universal approximation, the empirical results in the original paper demonstrate the great expressive power of KANs (Liu et al., 2025), motivating our empirical study.
@@ 4.2 The KAGCN Layer
GCN-based architectures have achieved great success in node classification tasks. While in our experiments we evaluate KAGIN on node classification datasets, the objective advantage of GCN over GIN on some of the datasets does not facilitate a fair estimation of KANs’ potential in this context. To this end, we also propose a variant of the GCN model. Specifically, we substitute the linear transformation of the standard GCN (Kipf and Welling, 2017) model with a single KAN layer (defined in Equation 8) to obtain the KAGCN layer:
@@
4.3 The KAGAT Layer
Finally, we substitute the linear transformation of the standard graph attention network (GAT)  (Veličković et al., 2017) model with a single KAN layer (defined in Equation 8) to obtain the KAGAT layer:
$\mathbf{h}^{(\ell)}_{v}=\sum_{u\in\mathcal{N}(v)\cup\{v\}}\alpha^{(\ell)}_{v,u}\Phi^{(\ell)}\left(\mathbf{h}_{u}^{(\ell-1)}\right)$ (11)
where $\Phi^{(\ell)}$ denotes a single KAN layer and
$\alpha_{uv}^{(\ell)}=\frac{\exp\left(\texttt{LeakyReLU}\left(\left[\Phi^{(\ell)}\left(h^{(\ell-1)}_{u}\right)+\Phi^{(\ell)}\left(h^{(\ell-1)}_{v}\right)\right]\right)\right)}{\sum_{u\in\mathcal{N}(v)\cup\{v\}}\left(\exp\left(\texttt{LeakyReLU}\left(\left[\Phi^{(\ell)}\left(h^{(\ell-1)}_{u}\right)+\Phi^{(\ell)}\left(h^{(\ell-1)}_{v}\right)\right]\right)\right)\right)}$ (12)
@@ 5 Empirical Evaluation
In this section, we compare all three KAGNN models against MLP-based GNNs on the following tasks: node classification, link prediction, graph classification, and graph regression. All models are implemented with PyTorch (Paszke et al., 2019). For KAN layers, we rely on publicly available implementations for B-splines  [^1] https://github.com/Blealtan/efficient-kan and for RBF  [^2] https://github.com/ZiyaoLi/fast-kan. For a given task, all models have consistent architectures, differing as little as possible. All GAT-based models use 4 heads.
@@ 5.1 Node classification
To evaluate the performance of GNNs with KAN layers in the context of node classification, we use $7$ well-known datasets of varying sizes and types, including homophilic (Cora, Citeseer (Kipf and Welling, 2017) and Ogbn-arxiv (Hu et al., 2020)) and heterophilic (Cornell, Texas, Wisconsin, Actor) networks (Zhu et al., 2021). The homophilic networks are already split into training, validation, and test sets, while the heterophilic datasets are accompanied by fixed $10$-fold cross-validation indices. Thus, we clone the split for homophilic datasets $10$ times in order to have a similar number of experiments across all datasets.
@@
We perform two types of experiments. First, we compare our KAGCN model on the homophilic datasets to the GCN proposed in the state of the art approach in (Luo et al., 2024). The results are shown in Table 1. One can observe that the classical GCN outperforms the KAN-based model in 2 out of 3 datasets, albeit with a narrow difference and confidence overlaps in all of the comparisons. Note that the current setting is built to favor the classical GNNs, as it relies on techniques such as dropout, jumping knowledge, batch normalization and preliminary linear transformations. Hence, although the MLP-based GCN enjoys a plethora of assistive techniques examined in the past decades, the KAN-based counterpart exhibits a close performance. We should note here that these assistive techniques have not been developed yet for KAN-based architectures, and our results aim at motivating the development of such techniques.
Table 1: Average classification accuracy ($\pm$ standard deviation) and model parameters based on Luo et al.(Luo et al., 2024) using GCN-based models on the Cora, CiteSeer, and Arxiv datasets.
@@
Our second approach compares the vanilla versions of the neural architectures when we replace the MLP with KAN, as delineated in sections 4. In order to clarify the effect of KAN on the model’s performance, we only use dropout and batchnorm as additional techniques. Indeed, we aspire to compare the classical MLP-based GNNs as if they were in the same nascent stage as their KAN-based counterparts.
In order for all models to exploit the same quantity of information, we fix the number of message-passing layer dataset-wise, with $2$ for Cora and CiteSeer, $3$ for Texas, Cornell, Wisconsin, and Ogbn-arxiv and $4$ for Actor. All attention-based models used 4 heads.
For every dataset and model, we tune the values of the hyperparameters using the Optuna package (Akiba et al., 2019) with $100$ trials (parameterizations), a TPE Sampler and set early stopping patience to $50$. For each parameterization, and for each split, we train $3$ models in order to mitigate randomness. We then select the parameterization with lowest validation error across all splits and models.
We use a different hyperparameter range for MLP-based models, B-Splines-based models and RBF-based models (see Table 8). Indeed, for the same number of ”neurons”, a KAN-based layer will have a much larger number of parameters than an MLP-based layer (see Appendix C.1 for examples). In order to avoid giving an unfair advantage to KAN-based layers, we thus allowed the MLP layers to be larger.
Table 2: Average node classification accuracies ($\pm$ standard deviation).
@@
GCN 74.01 $\pm$ 0.81 57.73 $\pm$ 1.21 53.21 $\pm$ 0.13 66.13 $\pm$ 6.09 72.34 $\pm$ 5.92 78.63 $\pm$ 5.43 32.81 $\pm$ 1.11
BS-KAGCN 70.46 $\pm$ 2.00 56.45 $\pm$ 1.62 53.76 $\pm$ 0.31 67.75 $\pm$ 6.54 72.16 $\pm$ 5.68 78.63 $\pm$ 5.26 34.45 $\pm$ 1.12
RBF-KAGCN 48.06 $\pm$ 1.40 47.55 $\pm$ 1.43 53.65 $\pm$ 0.20 67.75 $\pm$ 7.24 71.62 $\pm$ 4.69 77.45 $\pm$ 4.11 34.77 $\pm$ 1.20
GAT 76.51 $\pm$ 0.79 60.91 $\pm$ 1.04 53.68 $\pm$ 0.12 66.85 $\pm$ 5.59 69.01 $\pm$ 6.14 78.17 $\pm$ 5.70 32.51 $\pm$ 1.17
BS-KAGAT 78.28 $\pm$ 0.83 56.87 $\pm$ 1.67 55.17 $\pm$ 0.12 67.66 $\pm$ 6.94 73.78 $\pm$ 6.39 78.43 $\pm$ 5.47 31.66 $\pm$ 1.23
RBF-KAGAT 72.62 $\pm$ 1.64 48.65 $\pm$ 2.39 54.24 $\pm$ 0.25 64.14 $\pm$ 6.14 71.71 $\pm$ 6.49 76.34 $\pm$ 4.34 34.04 $\pm$ 0.94
GIN 70.85 $\pm$ 1.75 47.98 $\pm$ 0.91 53.58 $\pm$ 0.74 69.10 $\pm$ 4.85 73.51 $\pm$ 5.87 78.24 $\pm$ 4.76 33.46 $\pm$ 1.38
BS-KAGIN 65.78 $\pm$ 1.68 55.77 $\pm$ 1.51 53.05 $\pm$ 0.55 69.64 $\pm$ 7.40 72.97 $\pm$ 5.63 78.63 $\pm$ 5.16 34.26 $\pm$ 1.06
RBF-KAGIN 72.77 $\pm$ 1.35 46.87 $\pm$ 1.93 53.56 $\pm$ 0.37 67.30 $\pm$ 6.44 73.60 $\pm$ 5.05 76.73 $\pm$ 4.88 34.37 $\pm$ 0.95
@@ Results.
The results are given in Table 2. We see that a B-splines-based variant outperforms other models on $5/7$ datasets, and both other architecture win on one dataset each. Nonetheless, these are often by small margins. We also notice poor performance of RBF-KAGCN on Cora, which we attribute to overfitting. Overall,there seems to be a positive impact of using KAN-based architectures.
5.2 Link Prediction
Datasets and Experimental Setup.
In this setting, we focus on the task of link prediction. We use two datasets, Cora and CiteSeer, following the implementation and protocol provided in (Li et al., 2023; Mao et al., 2024). We performed a grid-search over learning rate, hidden dimension, number of hidden layers, grid size and spline order, each time training models on 10 splits. We then selected the model with lowest validation loss to obtain the best configuration, and reported its test ROC-AUC (over all 10 splits).
Table 3: AUC for link prediction of the KAGIN and GIN models on the link-prediction datasets.
Cora CiteSeer Cora CiteSeer Cora CiteSeer
GCN 95.36 $\pm$ 0.23 96.72 $\pm$ 0.44 GAT 94.02 $\pm$ 0.48 96.10 $\pm$ 0.49 GIN 94.97 $\pm$ 0.40 95.37 $\pm$ 0.23
BS-KAGCN 94.37 $\pm$ 0.46 94.91 $\pm$ 0.62 BS-KAGAT 94.71 $\pm$ 0.65 95.40 $\pm$ 0.33 BS-KAGIN 94.64 $\pm$ 0.44 95.75 $\pm$ 0.39
RBF-KAGCN 94.87 $\pm$ 0.45 95.44 $\pm$ 0.67 RBF-KAGAT 94.96 $\pm$ 0.42 90.09 $\pm$ 14.1 RBF-KAGIN 95.07 $\pm$ 0.27 97.07 $\pm$ 0.40
Results.
Table 3 displays the results on the link prediction task. We see that GCN is the best performer on Cora, and RBF-KAGIN is the best on CiteSeer, albeit with often close scores among models. We also notice some instability of RBF-KAGAT on CiteSeer.
@@ 5.3 Graph Classification
In this set of experiments, we evaluate the KAGNN models on standard graph classification benchmark datasets (Morris et al., 2020b). We experiment with the following $7$ datasets: MUTAG, DD, NCI1, PROTEINS, ENZYMES, IMDB-B, IMDB-M.
Experimental setup. We follow the experimental protocol proposed in (Errica et al., 2020). We kept our architecture simple: first, a number of message-passing layers extract node representation, whose model is either an MLP (resp. a KAN). We then use pooling (sum-pooling for GINs and GATs, mean-pooling for GCNs) to obtain graph representations. Finally, an MLP (resp. a KAN) with same architecture as those in the message-passing layer, is used as a classifier. A softmax is used for predicting class probabilities. Thus, we perform $10$-fold cross-validation, while within each fold a model is selected based on a $90\%/10\%$ split of the training set. We use the pre-defined splits provided in (Errica et al., 2020). We use the Optuna package to select the model that achieves the lowest validation cross-entropy loss over $100$ iterations. For fair comparison, we fix the number of message-passing layers of all architectures for each dataset. On MUTAG, PROTEINS, IMDB-B and IMDB-M, we set the number of layers to $2$. On DD, we set it to $3$. On ENZYMES, we set it to $4$ and finally, on NCI1, we set it to $5$. We train each model for $1,000$ epochs (early stopping with a patience of $20$ epochs) by minimizing the cross entropy loss. We use the Adam optimizer for model training (Kingma and Ba, 2015). We apply batch normalization (Ioffe and Szegedy, 2015) and dropout (Srivastava et al., 2014) to the output of each message-passing layer. The ranges for hyperparameters search are given in table 11. Once the best hyperparameters are found for a given split, we train $3$ different models on the training set of the split and evaluate them on the test set of the split. This yields $3$ test accuracies for this split, and we compute their average to obtain the performance for this split.
Table 4: Average graph classification accuracy ($\pm$ standard deviation).
@@
GCN 59.95 $\pm$ 8.98 65.87 $\pm$ 6.33 60.10 $\pm$ 4.63 63.10 $\pm$ 3.75 20.56 $\pm$ 3.70 68.40 $\pm$ 4.01 48.56 $\pm$ 3.71
BS-KAGCN 71.70 $\pm$ 7.54 72.74 $\pm$ 4.16 73.97 $\pm$ 4.01 75.68 $\pm$ 3.49 59.06 $\pm$ 4.48 72.87 $\pm$ 5.12 50.04 $\pm$ 3.99
RBF-KAGCN 77.86 $\pm$ 9.45 66.78 $\pm$ 5.91 73.07 $\pm$ 2.90 73.67 $\pm$ 3.99 56.83 $\pm$ 8.37 73.13 $\pm$ 4.5 49.84 $\pm$ 3.78
GAT 57.51 $\pm$ 12.3 71.92 $\pm$ 5.88 60.42 $\pm$ 8.21 59.78 $\pm$ 0.87 18.17 $\pm$ 1.34 67.37 $\pm$ 3.83 48.24 $\pm$ 5.02
BS-KAGAT 82.66 $\pm$ 4.78 78.27 $\pm$ 2.37 76.46 $\pm$ 3.66 72.75 $\pm$ 2.18 51.33 $\pm$ 12.9 72.00 $\pm$ 3.80 50.09 $\pm$ 4.34
RBF-KAGAT 75.88 $\pm$ 12.1 73.42 $\pm$ 4.76 72.63 $\pm$ 4.65 71.93 $\pm$ 4.23 49.78 $\pm$ 7.65 71.00 $\pm$ 5.04 48.13 $\pm$ 4.15
GIN 86.16 $\pm$ 5.96 74.78 $\pm$ 3.82 79.32 $\pm$ 1.63 70.66 $\pm$ 3.60 54.72 $\pm$ 6.55 73.37 $\pm$ 2.90 50.36 $\pm$ 4.10
BS-KAGIN 85.79 $\pm$ 8.06 76.06 $\pm$ 5.10 80.34 $\pm$ 1.46 73.37 $\pm$ 2.81 44.50 $\pm$ 8.53 73.13 $\pm$ 4.46 50.33 $\pm$ 3.87
RBF-KAGIN 85.10 $\pm$ 5.60 75.92 $\pm$ 3.49 79.19 $\pm$ 1.81 74.33 $\pm$ 3.52 44.78 $\pm$ 9.58 69.93 $\pm$ 3.84 49.22 $\pm$ 4.41